Operations Report – February 18th

Crew 222 Operations Report 02/18/2020

SOL: 2

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator: yes
Hours run: 9 hours
From what time last night: 9pm
To what time this morning: 6am
List any additional daytime hours when the generator was run: N/A

Solar—SOC 93% (Before generator is run at night)

Diesel Reading – 78%

Station Propane Reading – 39% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149317.02 units

Water (static tank) – 427 gallons

Static to Loft Pump used – yes

Water in GreenHab – 127 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: yes
Hours: 143.8
Beginning charge: 100
Ending charge: 99
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: no
Hours: 131.7
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.0
Beginning charge: 100
Ending charge: 92
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: N/A
Reason for use: N/A
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 0
Notes on ATVs: the second one from the hab doesn’t have anymore battery

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the EVA. All ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: Tried to fix 2 radios with two batteries found in the RAM but there is no back cover so we can’t charge them.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

EVA Report – February 18th

Crew 222 EVA Report 18-Feb-2020
EVA # 2
Author: Nicolas Bourdeaud
Purpose of EVA: Install the electric field sensor (named MegaARES). Start time: 9:26am
End time: 11:01am
After the first steps on Mars for three of us, we checked all rovers status and the static tank level.
Then we decided to check the ATV’s batteries. Three of them started more or less easily. After 5 minutes, as we did not succeed in switching on the last one, I decided to give up, and to let this issue for the engineer’s EVA of tonight.
Then we moved to the place we needed to set the experiment, the same we chose yesterday, for the weather station and the LOAC.
One team began by changing the LOAC’s battery.
The installation of the MegaARES has been successfully carried out. We came back earlier than what we had planned.

Destination: plane area approximately 500m away from the Hab, along Cow Dung Road (same than yesterday)
Coordinates (use UTM NAD27 CONUS): 12S 0518494 4250679
Participants: Nicolas Bourdeaud (HSO), Benjamin Auzou (CO), Marie Bochard (Journalist), Eishi Kim (AST)
Road(s) and routes per MDRS Map: Cow Dung Road
Mode of travel: Driving (needed for material transport)

Astronomy Report – February 18th

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/18/2020


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: In addition to the still active observations from yesterday (NGC 5248, NGC 4038), 7 new targets are to be imaged: NGC 1961, NGC 4490, NGC 4450, NGC 6503, NGC 3368, NGC 3115, NGC 598.

Images submitted with this report: The first Standard image of NGC 4254 is attached separately in today’s Photo report. It is to be compared to the Reference image taken prior to the mission.

Problems Encountered: None


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Greenhab Report – February 18th

Crew 222 GreenHab Report – 18/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 20°C

Hours of supplemental light: NA

Daily water usage for crops: 11,6 gallons

Daily water usage for research and/or other purposes: 2 Liters

Water in blue Tank: 127 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative: Today, I watered plants after being the habcom of the second EVA. After the lunch, I went to take care of the spirulina and separated some plants( salad and carrots) to let them grow more. Then I helped Nina, our crew scientist, to plant in moss her seeds that we will next deplace in the
culture tower that was installed today in the science dome. I went finally in the evening to watered a second time the plants and took some sage for our meal.

Harvest: 4g of sage

Support/supplies needed:

Journalist Report – February 18th

Here is the report of the journalist who was in the EVA today !

Crew 222 Journalist Report 18Feb2020

Author: Marie Bochard, Crew Journalist

Sol 2

Title: Head in the stars!

Today in EVA, we were tasked with setting up the experiment MegaARES, a giant aluminium sphere that measures the magnetic field of the planet. When we practiced putting it up before leaving, it did not take too long, but here it took us two whole hours! It’s hard to imagine how difficult it is moving in a spacesuit. Every small task becomes difficult. Just getting in the rover seat was tough, and trying to get back up after sitting down was even harder. This is in part due to the large backpack we have to carry around, since it is quite heavy and impairs our movements, and since this was our first EVA we kept bumping into each other lightly. Thankfully we quickly got the hang of it, and the suit no longer bothered us.

Upon our return, we could smell the delicious smell of a savory cake waiting for us. Martian food, when cooked to perfection, can rival with Earth cuisine! Especially when seasoned with GreenHab-grown greens that our GreenHab officer harvested yesterday.

Here, far from civilization, far from any man-made light, the night sky is beautiful. Today, Eishi showed us the pictures of a galaxy that he took last night. As the Astronomer, he has access to three telescopes here in the MDRS.

The first one is for finding objects in the night sky, such as galaxies and stars. It is fully automated and it can follow the objects that the Astronomer wishes to study during the night. It then activates the second telescope that takes astrophotography, and in the morning, the Astronomer has access to all the images taken by the telescope.

The third one is a manual solar telescope that is used to follow the Sun’s activity during the day. With it, we can see clearly the dark spots and solar eruptions. This will be crucial on Mars, because without a magnetic field all the radiation from the Sun directly hits the surface of the planet, and this can be fatal to Marsonauts who will live there. The Sun isn’t a still ball of plasma, it has phases where its activity increases or decreases, and following these cycles is important to ensure we can predict future activity and prevent any danger. Solar flares can be dangerous for the Earth too, as it interferes with electronic instruments! Even if it is millions of kilometers away, it has the power to make and destroy.

Sol Summary – February 18th

Crew 222 Sol Summary Report 18-02-2020
Summary Title: Get to know Mars better
Author’s name: Benjamin Auzou
Mission Status: Mission planning followed, experiments running Sol Activity Summary:
In the morning, the crew woke up with an other matinal workout session, that created muscle pains to some members. Then four crewmember went on EVA to install a field experiment next to the experiment installed yesterday.
At lunch, the crewmembers that stayed in the Hab prepared a salted cake and a delicious soup.
In the afternoon, crewmebers worked on their experiments in order to start getting results as soon as possible. Also we all passed a human factors experiment, directed by Eishi.
We finished the day with a relaxation session directed by our HSO, a full body scan. The reactions were very different between the members of the crew.
So we started to know Mars better with the experiments, but also by finding habits here.

Look Ahead Plan:
Tomorrow, an EVA team will make some maintenance work on the two experiments and continue with an exploration of the area north of the Hab. We hope to get the water monitoring working as soon as tomorrow evening.

Anomalies in work: none
Weather: some clouds but clear in general
Crew Physical Status: some muscle pains from sport sessions
EVA: MegaARES installation success, all members of the crew have performed one EVA now
Reports to be filed: Journalist report, GreenHab Report, Astronomy Report, EVA Report, EVA Request, HSO report
Support Requested: propane tank refilling, trashes in the engineering airlock

Copyright © The Mars Society. All rights reserved. | Main Site